OpenStudy (kittiwitti1):

A sled of mass 50 kg is pulled along snow-covered, flat ground. The static friction coefficient is 0.30, and the sliding friction coefficient is 0.10. --- a) What does the sled weigh? b) What force will be needed to start the sled moving? c) What force is needed to keep the sled moving at a constant velocity? d) Once moving, what total force must be applied to the sled to accelerate it 3.0 m/s/s? --- Answers: a. 490 N b. 150 N c. 49 N d. 200 N

OpenStudy (kittiwitti1):

Firstly, I need to make sure the coefficient variables are right. Static friction:\[\mu _{s}\] Sliding friction (I take this means kinematic?):\[\mu _{k}\]or\[f\]

OpenStudy (anonymous):

phew. okay, sleds. and yeah, sliding friction would be kinetic friction.

OpenStudy (anonymous):

coefficient of kinetic friction is \[\mu_k\]

OpenStudy (kittiwitti1):

Okay, gotcha. Is this right?\[\mu _{k} < \mu _{s}\]

OpenStudy (anonymous):

yah

OpenStudy (kittiwitti1):

... Now I'm lost...

OpenStudy (anonymous):

^_^ So first, what does the sled weigh, and what is the summation of forces for the first part (static friction)?

OpenStudy (kittiwitti1):

Part A\[w = mg = 50 \times 9.8 = 490 N\] Part B\[F = \mu N\]... I don't know what to do now..

OpenStudy (anonymous):

That second part is the summation of the forces in the x-direction; and this is for part B, so make sure to label that mu as mu s. For an object on a horizontal surface, what is its normal force?

OpenStudy (kittiwitti1):

Part B\[F = \mu _{s}N = 0.30N\]

OpenStudy (kittiwitti1):

I don't know either F or N..? o.o

OpenStudy (anonymous):

for an object on a horizontal sutface, the Normal force is equal to its weight; you know that the force it takes to start the object moving is when \[F≥\mu_sN\] So yeah, just like you wrote \[ F≥(.3)N\]

OpenStudy (kittiwitti1):

... GRAH. We did NOT learn this like that = n =;;

OpenStudy (anonymous):

just use equals if it makes more sense ^_^

OpenStudy (anonymous):

you're looking for the minimum force it would take, so it's an "equals' anyway

OpenStudy (kittiwitti1):

Well it's in the notes ... but used this for in-class problems:\[F _{A} = f = \mu N\]

OpenStudy (kittiwitti1):

We also used the other one with an = sign.

OpenStudy (anonymous):

whatever makes the most sense ^^

OpenStudy (kittiwitti1):

But when I learned it in the NOTES:\[-F _{f} \le \mu _{s}N\]

OpenStudy (kittiwitti1):

Or Fn as they say, but same thing right?

OpenStudy (anonymous):

\[F_n=N\] yah And yeah, the force of static friction will be less than or equal to mu s N

OpenStudy (kittiwitti1):

What about this?

OpenStudy (anonymous):

you mean why did I write greater than?

OpenStudy (kittiwitti1):

\[\Sigma F_{y} = N - W = ma\]ma cancels out to 0\[\Sigma F _{y} = N - W = 0\] W = 490 (leaving N out because there is a variable N)\[N = W = -490\]

OpenStudy (kittiwitti1):

Sorry I should have typed that all together.

OpenStudy (anonymous):

N=+490 N

OpenStudy (kittiwitti1):

Yeah sorry I messed that up lolol

OpenStudy (anonymous):

then you is correct in what you wrote ^^

OpenStudy (kittiwitti1):

WOOT! But how do I find Fa?

OpenStudy (kittiwitti1):

OWAIT!\[\mu _{k} = \frac{ F _{A} }{ N }\]\[0.10 = \frac{ F _{A} }{ 490 }\]\[F _{A} = 49\] I feel stupid now = n =;;

OpenStudy (anonymous):

but you figured it out - you should feel smart! ^_^ If you set that up with a force summation you get the same thing. When the sled is moving, it's moving at a constant velocity, so a=0 as well \[ \sum F_x = F_A - F_{kinetic friction} = 0 \ \ \longrightarrow \ \ F_A = \mu_kN\]

OpenStudy (anonymous):

Got the last bit figured out?

OpenStudy (anonymous):

HALLO?? @kittiwitti1

OpenStudy (kittiwitti1):

Oh um ehh I was afk? xD

OpenStudy (anonymous):

figured you had fallen asleep at the keyboard ^_^

OpenStudy (kittiwitti1):

not that =_=" I actually took a nap beforehand

OpenStudy (kittiwitti1):

Wait - why do you even need the first part of the equation?

OpenStudy (anonymous):

you used the definition of the coefficient of kinetic friction - the way I did it derived it from newton's second law. They're equivalent ^_^ Is that what you meant?

OpenStudy (kittiwitti1):

No, I meant why do you even have to type it out...

OpenStudy (kittiwitti1):

**Write it out

OpenStudy (anonymous):

habit? I don't get what you mean...

OpenStudy (anonymous):

oh! You were using FA to be the actual force of kinetic friction! I get it. I thought FA was the force being applied to the sled. \[F_{push} - F_{k \ friction} = 0 \longrightarrow\quad F_{push} = F_{k \ friction} \longrightarrow \quad F_{push} =\mu_kN\]

OpenStudy (kittiwitti1):

Wait WHAT?! NO!! o_o I'm using Fa for applied force!!

OpenStudy (anonymous):

Then I really don't understand your question... :/

OpenStudy (kittiwitti1):

@ _ @ I don't get it anymore either

OpenStudy (anonymous):

you have the right answer...

OpenStudy (kittiwitti1):

Is the front part of the equation necessary to write out*

OpenStudy (anonymous):

If you say it's newton's second law yes; if you say it's the third, no, since there are only two forces in this system

OpenStudy (kittiwitti1):

We learned to use this formula:\[F = \mu _{k}N\]

OpenStudy (anonymous):

That's the definition for the force of kinetic friction, yah, and setting it equal to "force applied" works in this instance because it's the only other force in the problem

OpenStudy (kittiwitti1):

We didn't learn the first part...

OpenStudy (kittiwitti1):

I mean, F kin. friction? What is that o_o

OpenStudy (anonymous):

Everything you do until forever in mechanics always relates back to Newton's Second law. In this problem, the force opposing the motion of the sled is friction, so you can always set up a force diagram and sum the forces like what I typed out up there.

OpenStudy (kittiwitti1):

Is it okay to skip it?

OpenStudy (kittiwitti1):

WAIT do you mean Fa - f = ma?

OpenStudy (anonymous):

yesyesyesyesyes

OpenStudy (kittiwitti1):

OHHHH lols wow I'm so stupid xD

OpenStudy (anonymous):

\[ \sum F_x = F_{applied} - F_{friction} = \cancel{ma}^0\]

OpenStudy (kittiwitti1):

How did you cross that out o.o

OpenStudy (anonymous):

\cancel{ma}^0

OpenStudy (kittiwitti1):

no on the console what's the button for it

OpenStudy (anonymous):

there isn't one - I don't use the console

OpenStudy (kittiwitti1):

WOW. tech genius lol

OpenStudy (anonymous):

\sum F_x = F_{applied}-F_{friction} = \cancel{ma}^0 waaaay faster

OpenStudy (anonymous):

oh, I meant I don't use the buttons

OpenStudy (kittiwitti1):

Exactly, so you figured that out on your own! :p \[Sum F_x = F_a - f = \cancel{ma}^0\] I did "[Sum F_x = F_a - f = \cancel{ma}^0"

OpenStudy (kittiwitti1):

Thanks for the tip~

OpenStudy (anonymous):

welcome - \[ \LaTeX\] is wonderful!! you should learn it

OpenStudy (anonymous):

every operation has to have \ before it, like \sum \[\sum\] then \[ at the beginning of the phrase and

OpenStudy (anonymous):

\] at the end

OpenStudy (anonymous):

\[ \text{ \[ \sum F = ...... = \cancel{ma}^0\] }\]

OpenStudy (anonymous):

needs \

OpenStudy (anonymous):

the symbol right below the delete key

OpenStudy (anonymous):

three right of "P"

OpenStudy (anonymous):

in any case, the last part isn't too bad \[ \sum F_x = F_{applied} - F_{k \ friction} = ma \quad ; \quad a=3m/s^2 \\ \ \\ \ \\ \ \qquad F_{applied} = ma + F_{k \ friction} = ma + \mu_k N\]

OpenStudy (kittiwitti1):

O_O wat

OpenStudy (anonymous):

another Newton's second law

OpenStudy (anonymous):

The net force on the sled is \[ma\] , and the two forces that are acting on it are \[F_{applied}\] and \[F_{kinetic friction}\]

OpenStudy (kittiwitti1):

ok I'm really confused now, are we doing c) or d)? = n =;;

OpenStudy (anonymous):

So to have the sled accelerate at 3m/s^2, the force applied to the sled has to overcome the force of kinetic friction, and also accelerate the sled at 3 m/s^2 - therefor, the force applied must equal the force of kinetic friction plus the mass of the sled times its acceleration (3m/s^2)

OpenStudy (anonymous):

Oh, this is for the last part. I'm sorry. I thought you had already done C...

OpenStudy (kittiwitti1):

I'm not done with c)

OpenStudy (rajat97):

for the part a of the question you were right for the second part, you should just find the normal force between the block and the ground and then multiply it with the coefficient of static froction to get the limiting condition this means that if you apply force of magnitude greater than this static friction, the sledge will start moving next for part c as you have dine it before, that total forces on the sledge should be zero, so it will move with zero acceleration so, it will have constant velocity. but this time, the block is moving so, you should multiply the coeff. of kinetic friction(specified as sliding friction) with the normal force to get the frictional force next for the part d, f=ma f=50x3=150 and a force of 49 n is already acting on the block to make it move with a constant velocity. thus by adding both the forces i.e. fand 49n, we will get the total force=199n nearly equal to 200n

OpenStudy (anonymous):

You wrote the answer down up there \[F_A = \mu_kN = (.10)(490N) \\ \\ \ F_A = 49N\]

OpenStudy (anonymous):

A sled of mass 50 kg is pulled along snow-covered, flat ground. The static friction coefficient is 0.30, and the sliding friction coefficient is 0.10. --- a) What does the sled weigh? b) What force will be needed to start the sled moving? c) What force is needed to keep the sled moving at a constant velocity? d) Once moving, what total force must be applied to the sled to accelerate it 3.0 m/s/s? --- Answers: a. 490 N b. 150 N c. 49 N d. 200 N

OpenStudy (kittiwitti1):

.... nvm

OpenStudy (kittiwitti1):

I misread...

OpenStudy (kittiwitti1):

Wait, then how to get b)?

OpenStudy (anonymous):

\[F_A=\mu_sN\] \[F_A=(.3)(490N)\]

OpenStudy (kittiwitti1):

NEVERMIND! I got it.

OpenStudy (anonymous):

^_^

OpenStudy (kittiwitti1):

ok what's d)?

OpenStudy (kittiwitti1):

from @rajat97 next for the part d, f=ma f=50x3=150 and a force of 49 n is already acting on the block to make it move with a constant velocity. thus by adding both the forces i.e. fand 49n, we will get the total force=199n nearly equal to 200n

OpenStudy (kittiwitti1):

idk, it's confusing

OpenStudy (anonymous):

Write out the forces acting on the sled

OpenStudy (kittiwitti1):

idk my homework is due in less than an hour and I have 5 problems left -_-;;

OpenStudy (kittiwitti1):

wait Fa(k) + Fa(s)?

OpenStudy (anonymous):

I don't know what that means